OpenStudy (anonymous):

Find the intersection point(s) for the following functions: a(x) = 2x2 + 12x - 17 b(x) = x2 + 12x - 1

OpenStudy (anonymous):

equate the two equations then solve for x and then use each x to substitute to solve for a(x) and b(x)

OpenStudy (anonymous):

um how do i evaluated the equation?

OpenStudy (anonymous):

the same way as the previous question you asked it is the same method. follow that

OpenStudy (anonymous):

ok. thank you

OpenStudy (anonymous):

so it is going to be 2x^2+12x-17=x^2+12x-1 then solve for x the answers that you get you substiute to get the y co-ordinates

OpenStudy (anonymous):

i get it. well a bit. Find the intersection point(s) for the following functions: a(x) = 4x2 + 7x - 53 b(x) = 3x2 + 7x - 4 for these questions, i see some of the answer choice have 2 answers. i can only get one part of it. how do i get the second part?

OpenStudy (anonymous):

the answer i had was (7, 192) but theres two answer choice for that. one with that answer alone and another with another answer. (7, 192) and (-7, 94)

OpenStudy (anonymous):

oh nevermind. i got it.

OpenStudy (anonymous):

cool